# Joint Reconstruction of Spatially-Coherent and Realistic Clothed Humans and Objects from a Single Image

### v1

@@ Abstract
Recent advances in human shape learning have focused on achieving accurate human reconstruction from single-view images. However, in the real world, humans share space with other objects. Reconstructing images with humans and objects is challenging due to the occlusions and lack of 3D spatial awareness, which leads to depth ambiguity in the reconstruction. Existing methods in monocular human-object reconstruction fail to capture intricate details of clothed human bodies and object surfaces due to their template-based nature. In this paper, we jointly reconstruct clothed humans and objects in a spatially coherent manner from single-view images, while addressing human-object occlusions. A novel attention-based neural implicit model is proposed that leverages image pixel alignment to retrieve high-quality details, and incorporates semantic features extracted from the human-object pose to enable 3D spatial awareness. A generative diffusion model is used to handle human-object occlusions. For training and evaluation, we introduce a synthetic dataset with rendered scenes of inter-occluded 3D human scans and diverse objects. Extensive evaluation on both synthetic and real datasets demonstrates the superior quality of proposed human-object reconstructions over competitive methods.
@@ 1 Introduction
In recent years, there has been a growing demand for digitizing 3D humans and objects for applications ranging from virtual reality, movie production, gaming and social telepresence. Research has moved towards obtaining high-quality surface details of either clothed humans [29, 41, 7, 22, 8] or textured objects [3, 45, 6, 33, 1] from single-view images. Since most existing methods focus solely on either human or object reconstruction, they perceive humans and objects in isolation, whereas real images generally contain both humans and objects in the same scene. Recent works [36, 47, 20, 38, 37] have reconstructed humans and objects jointly, but they rely on parametric template-based representations, which limit realism, failing to capture loose clothing, complex hairstyles, and the free-form shapes of objects. In this paper, for the first time, we jointly reconstruct realistic, non-parametric clothed humans and objects from single images, in a spatially coherent way, addressing the gap in the literature.
@@
Figure 1: SCHOR jointly reconstructs spatially coherent realistic clothed humans and objects from synthetic (a) and real (b) images by first handling occlusion with a conditioned generative model followed by attention-based neural implicit model estimation.
Single-view 3D reconstruction of human and object scenes is an inherently ill-posed problem. Multiple 3D spatial configurations of humans and objects can project to the same 2D image, leading to depth ambiguity and difficulty in estimating the relative size, shape and position of the human-object in the 3D space. Human-object occlusion further complicates the inference. Previous methods solve this by using parametric models [47] to enforce geometric and spatial constraints to predict the 3D spatial arrangement. However, they do not achieve realistic 3D human-object reconstruction, because this also requires recovering fine details in both visible and occluded regions, increasing the task complexity.
To address these challenges, we propose SCHOR, a novel framework that can jointly produce Spatially-Coherent realistic Human Object Reconstructions from a single image, as shown in Fig. 1. This is achieved by introducing a novel attention-based neural implicit model that estimates the implicit representation of both humans and objects. Unlike parametric representations, the implicit representation enhances the realism of the reconstruction because the proposed implicit model leverages local pixel-aligned features extracted from a full-body human image and a segmented object image. Because the human is likely occluded by the object, the full-body human image is created by inpainting the non-visible body regions of the segmented partial human image using a generative model, allowing the implicit model to reproduce details in the body regions occluded by the object. To ensure 3D spatial coherence, the neural implicit model uses self-attention to incorporate global context from the input human-object RGB image as well as semantic features from a human-object pose estimated from the image. As a result, SCHOR can jointly reason about the 3D human-object spatial configuration while also retrieving high-quality details in the reconstruction.
@@
Due to the lack of existing datasets with high-quality, ground-truth 3D surfaces for human-object scenes, we created synHOR, a synthetic dataset for Human Object Reconstruction to train and evaluate SCHOR. We simulate human-object scenes by randomly placing 3D human scans from the public THuman2.0 [44] dataset with selected 3D object meshes from BEHAVE [2] and HODome [46] datasets, in a variety of 3D spatial configurations. We also evaluate our model on the real-world BEHAVE [2] dataset and demonstrate superior performance of reconstructions against state-of-the-art methods. Our key contributions include:
@@
  • We introduce a novel framework to jointly reconstruct spatially coherent, realistic human-object shapes from single images. To the best of our knowledge, this is the first work towards achieving realistic non-parametric human-object reconstructions from single images.
  • We propose a novel attention-based neural implicit network that jointly reasons about the 3D human-object locations and retrieves realistic human surface details and object shapes.
@@
  • We demonstrate superior human-object reconstruction quality compared to the state-of-the-art methods, both quantitatively and qualitatively, on synthetic and real datasets.
@@ 2 Related works
3D human and object reconstruction: To reconstruct 3D human and object jointly, previous methods use parametric models to fit human and object meshes satisfying various human-object interaction constraints. PHOSA [47] and D3D-HOI [43] each proposed an optimization based framework with physical constraints on scale and predefined contact priors. Wang et al. [35] modeled 3D human-object interactions from an image using commonsense knowledge from large language models. Holistic++ [4] modeled fine-grained human-object relations in a scene and inferred using Markov chain Monte Carlo method. CHORE proposed the learning of neural-implicit functions and then fit parametric model to the learned representation. Vistracker [37], InterTrack [39] reconstruct single human-object from a video, by specifically modeling the temporal context. Recently, CONTHO [20] proposed a method to refine human-object reconstruction from an image by 3D guided contact estimation. ProciGen [38] designed a framework to synthetically generate a large diverse dataset, and proposed a Hierarchical Diffusion Model to reconstruct human and object. None of the current approaches focused on single-view human-object reconstruction are able to recover clothed human subjects at the same fidelity as SCHOR. Alternate methods [31, 9, 10, 46] that can reconstruct clothed human-object, either use sparse or multi-views or monocular RGBD image as inputs, thereby relaxing their constraints and are thus not comparable to our work. We address the problem of jointly reconstructing realistic non-parametric clothed humans and objects from single images.
Neural implicit model for single-view 3D: Early reconstruction methods cannot produce high-quality 3D shapes, which limits the realism of reconstructions. This arises from the discretized nature of traditional representations like voxels, meshes or point clouds. The ground-breaking introduction of the implicit representation for 3D reconstruction [19, 5, 21] leads deep learning approaches to adopt this continuous form of representation because it can represent fine-scale details on the reconstructed surfaces. These methods employ neural implicit models to estimate the implicit representation of 3D points in space.
Several works have since used neural implicit models for object reconstruction from single images. Early models estimate occupancy [5, 19] or signed distance fields [21, 42, 49] using multi-layer perceptrons (MLPs) that are conditioned on input features. More recent works modify this basic architecture to improve the object reconstruction by incorporating prior knowledge into implicit functions  [16, 3]; using monocular geometric cues  [45]; combining explicit templates with implicit representations [6, 33] or leveraging input global and local features  [14, 1].
Another line of neural implicit models focuses on reconstructing high-quality 3D human shapes from single images. PIFu [28] introduced the pixel-alignment implicit function, which enhanced the level of detail of the 3D human shape. Building on pixel alignment, several methods have been developed to further improve detail quality using normal maps [29, 40] or super-resolution shapes [22]; addressing depth ambiguity using parametric models [50, 7]; and achieving complete reconstructions with diffusion models [30, 12, 8] or by incorporating additional data such as depth maps [41, 23] or unconstrained images [24]
However, these approaches cannot jointly reconstruct both human and object shapes and are generally category-specific, with human-focused methods unable to reconstruct objects and object-focused methods unable to reconstruct human shapes. Our proposed approach enables spatially coherent and realistic joint reconstruction of clothed humans and objects from a single image.
3 Spatially-coherent, realistic 3D shapes of clothed humans and objects
Overview:
@@ 3 Spatially-coherent, realistic 3D shapes of clothed humans and objects
Figure 2: SCHOR overview: Given an input image of a human-object scene, we first use a generative model to inpaint occluded human body regions, guided by a mask of missing areas and the segmented input of the human. Next, the generated image, along with an estimated normal map, the input image, the segmented object image, and estimated pose parameters, are processed by an attention-based neural implicit model. This model jointly estimates the implicit representation of the human-object shape.
@@
Recent methods for modeling human-object interactions, rely on template-based parametric representations that lack high-quality details, such as clothing and wrinkles, reducing the realism of the reconstruction. Other works that aim to achieve high-quality 3D shapes are mostly category-specific (either object or human) and cannot jointly reconstruct human-object shapes. To address these challenges, we introduce a novel framework that takes an RGB image as input and achieves Spatial-Coherent, realistic clothed Humans and Objects Reconstructions (SCHOR). In the proposed pipeline, shown in Fig. 2, the image is first segmented to separate the human from the object, and pose parameters of the SMPL-H [27] model are estimated. Occlusion is then addressed using the generative power of an image-conditioned diffusion model, which inpaints occluded regions of the human body. Implicit representations of both the human $s_{h}$ and the object $s_{o}$ are then estimated using a novel attention-based neural implicit model that incorporates the input RGB image $I_{f}$, the generated full-body human image $I_{h}$, the object image $I_{o}$ and the estimated pose parameters. Compared to related works, our approach jointly reconstructs realistic clothed humans and objects while preserving 3D spatial coherence between them.
@@ 3.1 Inpainting of occluded human body regions
This work addresses cases in real images where the human is occluded by an object. When such occlusion happens, regions of the human body are missing in the input image, leaving the neural implicit model without the information needed to estimate these regions. We propose to leverage the generative capability of diffusion models to inpaint the occluded body regions, using the partial human image $I_{p}$ along with a mask of the occluded regions $M_{i}$. Given the input image $I_{f}$, we first apply semantic segmentation to separate the human from the object, while estimating SMPL-H [27] poses for both. Using the SMPL-H model, we then generate a mask of the human $M_{s}$, and we obtain the human-object intersection mask $M_{i}$ as $M_{i}=M_{s}-M_{p}$, where $M_{p}$ is the segmented partial human mask.
@@
Figure 3: The attention-based neural implicit model first extracts pixel-aligned features from the input images to capture local high-frequency details. It then uses a transformer encoder to merge these features, achieving 3D spatial awareness. Finally, the model estimates the implicit representations for both humans and objects.
@@ 3.2 Attention-based neural implicit model
The proposed implicit model comprises of multiple modules, as shown in Fig. 3
@@
Pixel-aligned feature extractors: Previous methods on 3D human reconstruction [28, 29, 22] demonstrate that projecting a 3D point $x\in\mathbb{R}^{3}$ in the embedded image feature space $\phi(I)$ extracted with a convolutional stacked hourglass network significantly increases the quality of 3D human shapes. SCHOR first extracts a feature embedding for each input image $\phi_{\{h,o,f\}}(I_{\{h,o,f\}})$. A first set of points $X_{h}$ is then projected onto $I_{f}$ and $I_{h}$ via perspective projection $\pi$ and linked to the corresponding features $\phi_{h}$ and $\phi_{f}$ to obtain pixel-alignment with the human: $\Phi_{\{h,f\}}^{h}=\phi_{\{h,f\}}(\pi(X_{h},I_{\{h,f\}}))$. Similarly, the pixel-aligned object features $\Phi_{\{o,f\}}^{o}=\phi_{\{o,f\}}(\pi(X_{o},I_{\{o,f\}}))$ are obtained by projecting a different set of points $X_{o}$ on $I_{o}$ and $I_{f}$ and indexing them with the corresponding features $\phi_{o}$ and $\phi_{f}$.
@@
Human-object pose prior features: To guide the 3D reconstruction, we use geometric priors from an initial SMPL-H and object mesh estimated from the input image. Following previous work [36, 37, 20], our method initializes the object pose from a given object template. To define the pose features, for a query point $x_{h}\in X_{h}$, we look for the closest point $x_{h}^{*}$ on SMPL-H, i.e. $x_{h}^{*}=\arg\min_{x_{h}^{p}}||x_{h}-x_{h}^{p}||$ where $x_{h}^{p}$ are points on the SMPL-H mesh. Our human pose prior features $\sigma_{h}(x_{h})=[d_{h},v_{h},z_{h}]$ then constitutes three elements: a signed distance value $d_{h}$ between $x_{h}$ and $x_{h}^{*}$, a visibility label $v_{h}\in\{1,0\}$ where $v_{h}$ indicates if $x_{h}^{*}$ is visible in the front image when SMPL-H and initial object mesh is projected together, and a relative depth-aware feature $z_{h}=(x_{h}^{z}-z_{c})$ where $z_{c}$ is the depth of the SMPL-H centre and $x_{h}^{z}$ is the depth of the query point $x_{h}$. Analogously, for a query point $x_{o}\in X_{o}$ and its closest point on the object mesh, $x_{o}^{*}$, the object pose prior features $\sigma_{o}(x_{o})=[d_{o},v_{o},z_{o}]$ also constitutes three elements: a signed distance value $d_{o}$ between $x_{o}$ and $x_{o}^{*}$, a visibility label $v_{o}\in\{1,0\}$ and a relative depth-aware feature $z_{o}=(x_{o}^{z}-z_{c})$ where $x_{o}^{z}$ is the depth of the query point $x_{o}$. The object position is also defined relative to the SMPL-H centre, to learn 3D spatial coherency with the human.
@@
Global spatial awareness via transformer encoders: as shown in Sec. 4.1, simply concatenating the features of the human $\Phi_{h}^{h}$ and object $\Phi_{o}^{o}$ with those of the input image $\Phi_{f}^{\{h,o\}}$ deteriorates network performance, making it less robust to noise of the input image. Additionally, the network’s ability to integrate and reason about features across the entire image is limited by the receptive field of convolutional layers. To achieve 3D spatially coherent reconstruction, we propose using two attention-based encoders, $A_{o}$ and $A_{h}$, to merge the feature extracted by the feature extractors module. An attention score is computed for the two paired images by assessing the compatibility between a query and its corresponding key, producing two feature representations: one for the human $\varphi_{h}=A_{h}(\Phi_{h}^{h},\Phi_{f}^{h},\sigma_{h}(x_{h}))$ and another for the object $\varphi_{o}=A_{o}(\Phi_{o}^{o},\Phi_{f}^{o},\sigma_{o}(x_{o}))$. Notice that the pose features $\sigma_{\{h,o\}}$ obtained in the previous module are also concatenated. Each feature $\varphi_{\{h,o\}}$ integrates the information from its corresponding image $I_{\{h,o\}}$, combined with the global context of the input image $I_{f}$. This enables a more comprehensive understanding of the scene by using the multi-head self-attention layers of the transformer. As a result, the human feature $\varphi_{h}$ contains detailed information from the human image, as well as contextual information of the scene from the input image. Similarly, the object feature $\varphi_{o}$ integrates scene-level information from $I_{f}$.

### v2

# Realistic Clothed Human and Object Joint Reconstruction from a Single Image

@@ Abstract
Recent approaches to jointly reconstruct 3D humans and objects from a single RGB image represent 3D shapes with template-based or coarse models, which fail to capture details of loose clothing on human bodies. In this paper, we introduce a novel implicit approach for jointly reconstructing realistic 3D clothed humans and objects from a monocular view. For the first time, we model both the human and the object with an implicit representation, allowing to capture more realistic details such as clothing. This task is extremely challenging due to human-object occlusions and the lack of 3D information in 2D images, often leading to poor detail reconstruction and depth ambiguity. To address these problems, we propose a novel attention-based neural implicit model that leverages image pixel alignment from both the input human-object image for a global understanding of the human-object scene and from local separate views of the human and object images to improve realism with, for example, clothing details. Additionally, the network is conditioned on semantic features derived from an estimated human-object pose prior, which provides 3D spatial information about the shared space of humans and objects. To handle human occlusion caused by objects, we use a generative diffusion model that inpaints the occluded regions, recovering otherwise lost details. For training and evaluation, we introduce a synthetic dataset featuring rendered scenes of inter-occluded 3D human scans and diverse objects. Extensive evaluation on both synthetic and real-world datasets demonstrates the superior quality of the proposed human-object reconstructions over competitive methods.
@@ 1 Introduction
Realistic, personalized human avatars that seamlessly coexist with objects will shape the future of movies, games, telepresence, and the metaverse. The joint reconstruction of clothed humans and objects will be key to this vision. Emphasis will be on achieving realism in the reconstructed shapes to accurately reflect real-world characteristics.
@@
Figure 1: ReCHOR jointly reconstructs realistic clothed humans and objects from synthetic (a) and real (b) images by first handling human occlusion with a conditioned generative model followed by attention-based neural implicit model estimation.
Motivated by this, we aim to jointly reconstruct realistic clothed humans and objects from a single-view human-object scene. This task presents significant challenges due to human-object occlusions and unknown camera parameters, which make it difficult to accurately infer the 3D spatial configuration (depth, scale, pose) as well as the realistic shape for reconstruction. Existing methods  [34, 45, 18, 36, 35] that reconstruct 3D humans and objects from a single RGB image focus mainly on optimizing the 3D spatial configuration, failing to capture realistic details such as clothing, hairstyles, and the free-form geometry of human-object shapes. They represent 3D shapes using either parametric, template-based, or coarse models, which constrain the surface geometry, thereby limiting realism. We therefore explore implicit representations [17, 5, 19], which, unlike parametric models, enable realistic reconstruction without constraining the topology. However, while implicit representations can model realistic shapes and poses, they are prone to depth ambiguity if the only input is a 2D image without any explicit 3D depth information.
To address these challenges, we propose ReCHOR, a novel framework for Realistic Clothed Human and Object joint Reconstruction. To obtain realistic surface details, ReCHOR incorporates a novel attention-based neural implicit model to estimate implicit representations of human-object shapes, assisted by a generative diffusion model that recovers details from regions of the human body occluded by the object. To resolve depth ambiguity, the neural implicit model is also conditioned on an estimated human-object pose prior, integrating 3D spatial information into the estimation process. Specifically, we first segment the input human-object image to obtain separate human and object images. The generative diffusion model then inpaints the body regions occluded by the object in the human image, generating a full-body human image. This image, along with the object image and additional inputs, form the ’local’ context which embeds local details information. The input human-object image serves as the ’global’ context, providing spatial cues between the human and the object. The neural implicit model uses pixel-aligned features from both image contexts, merging them through an attention-based architecture before estimating implicit representations, allowing the retrieval of realistic details while considering the contextual relationship between the human and the object.
Additionally, the neural implicit model is conditioned on semantic features derived from human-object pose priors, which are estimated using parametric human-object reconstruction methods [34, 45, 18, 36]. These methods enforce geometric and spatial constraints to optimize for the 3D location, depth and scale of the human and object, providing ReCHOR with 3D spatial information essential to address the problem of depth ambiguity. The proposed model then computes the implicit representations in the reference frame of the predicted 3D location prior. By effectively decoupling depth ambiguity and detail surface retrieval, ReCHOR enables more accurate reconstruction of realistic clothed humans and objects, as shown in Fig. 2.
@@
Due to the lack of datasets with high-quality 3D ground-truth human-object scenes, we create synHOR, a synthetic dataset for Human Object Reconstruction to train and evaluate ReCHOR. We generate several 3D spatial configurations of human-object scenes by randomly placing 3D human scans from THuman2.0 [42] with selected object meshes from BEHAVE [2] and HODome [44]. We also evaluate our model on the real-world BEHAVE [2] dataset and demonstrate superior performance of reconstructions against state-of-the-art methods. Our key contributions include:
@@
  • A novel framework to jointly reconstruct realistic clothed humans and objects from single images. This is the first work that represents realistic human details in the joint reconstruction of a non-parametric human-object shape.
  • A novel attention-based neural implicit network to estimate the implicit representation of realistic clothed humans and objects. Pixel-aligned features are extracted from local and global views and then merged along with 3D spatial information via transformer encoders, capturing realistic details while learning contextual information across local and global scenes.
@@
  • We demonstrate superior reconstruction quality compared to the state-of-the-art methods, both quantitatively and qualitatively, on synthetic and real datasets.
@@ 2 Related works
3D human and object reconstruction: To reconstruct 3D human and object jointly, previous methods use parametric models to fit human and object meshes satisfying various constraints. PHOSA [45] and D3D-HOI [41] each proposed an optimization based framework with physical constraints on scale and predefined contact priors. Wang et al. [33] modeled 3D human-object shapes from an image using commonsense knowledge from large language models. Holistic++ [4] modeled fine-grained human-object relations in a scene using Markov chain Monte Carlo method. CHORE proposed to fit a parametric model to a learned neural-implicit functions. Vistracker [35], InterTrack [37] reconstruct human-object from a single video, by specifically modeling the temporal context. Recently, CONTHO [18] proposed a method to refine human-object reconstruction from an image by 3D guided contact estimation. ProciGen [36] proposed a Hierarchical Diffusion Model to reconstruct human and object. None of the current approaches for single-view human-object reconstruction can recover realistic clothed humans at the same fidelity as ReCHOR. Alternate methods [29, 9, 10, 44] that can reconstruct clothed human-object, either use sparse, multi-views or monocular RGBD image as inputs, thereby relaxing their constraints and are thus not comparable to our work.
Neural implicit model for single-view 3D reconstruction: Early reconstruction methods cannot produce realistic 3D shapes due to the discretized nature of traditional representations like voxels, meshes or point clouds. The introduction of the implicit representation for 3D reconstruction [17, 5, 19] has led deep learning approaches to adopt this continuous representation because it can represent fine details on the reconstructed shapes. These works estimate the implicit representation with neural implicit models. Several works have since used neural implicit models for object reconstruction from single images. Early models estimate occupancy [5, 17] or signed distance fields [19, 40, 47] using multi-layer perceptrons conditioned on input features. Recent works improve the object reconstruction by incorporating prior knowledge into implicit functions  [15, 3]; using monocular geometric cues  [43]; combining explicit templates with implicit representations [6, 31] or leveraging input global and local features  [14, 1]. Other implicit models focus on reconstructing high-quality 3D human shapes from single images. PIFu [26] introduced the pixel-aligned implicit function to retrieve detailed human shapes. Building on this, several works improved quality using normal maps [27, 38] or super-resolution shapes [20]; addressed depth ambiguity using parametric models [48, 7]; achieved complete reconstructions with diffusion models [28, 12, 8] or incorporating additional data as depth maps [39, 21] or unconstrained images [22]. These works cannot jointly reconstruct human-object shapes since they are category-specific, with human-focused methods unable to reconstruct objects and vice-versa. ReCHOR jointly reconstructs realistic clothed humans and objects from a single image.
3 Realistic 3D shapes of clothed humans and objects
@@
Figure 2: ReCHOR overview: Given an input image of a human-object scene, we first use a generative model to inpaint occluded human body regions, guided by a mask of missing areas and the segmented input of the human. Next, the generated image, along with an estimated normal map, the input image, the segmented object image, and estimated pose parameters, are processed by an attention-based neural implicit model. This model jointly estimates the implicit representation of the human-object shape.
@@
We introduce ReCHOR, a novel framework for Realistic Clothed Humans and Objects Reconstructions from a single RGB image depicting both the human and the object. In the proposed pipeline, shown in Fig. 2, the image is first segmented to separate the human from the object, and pose parameters of the SMPL-H [25] model and object are estimated. Regions of the human body that are occluded by the object in the input image are inpainted using the generative power of an image-conditioned diffusion model. Implicit representations of both the human $s_{h}$ and the object $s_{o}$ are then estimated using a novel attention-based neural implicit model that incorporates the input RGB image $I_{f}$, the generated full-body human image $I_{h}$, the object image $I_{o}$ and the estimated pose parameters. Compared to related works, our approach jointly reconstructs realistic clothed humans and objects while avoiding depth ambiguity between them.
@@ 3.1 Inpainting of occluded human body regions
This work addresses cases in images where an object occludes the human. When such occlusion happens, regions of the human body are missing in the input image, leaving the neural implicit model without the information needed to estimate these regions. We propose to leverage the generative capability of diffusion models to inpaint the occluded body regions, using the partial human image $I_{p}$ along with a mask of the occluded regions $M_{i}$. Given the input image $I_{f}$, we first apply semantic segmentation to separate the human from the object, while estimating the SMPL-H [25] and object pose. Using the SMPL-H model, we then generate a mask of the human $M_{s}$, and we obtain the human-object intersection mask $M_{i}$ as $M_{i}=M_{s}-M_{p}$, where $M_{p}$ is the segmented partial human mask.
@@
Figure 3: The attention-based neural implicit model first extracts pixel-aligned features from the input images to capture local details. It then uses a transformer encoder to merge these features, learning global and local contextual information about the scene. Finally, the model estimates the implicit representations for both humans and objects. Human-object pose priors provide 3D spatial information to address depth ambiguity.
@@ 3.2 Attention-based neural implicit model
The proposed implicit model comprises of multiple modules, as shown in Fig. 3.
@@
Pixel-aligned feature extractors: Previous methods on 3D human reconstruction [26, 27, 20] demonstrate that projecting a 3D point $x\in\mathbb{R}^{3}$ in the embedded image feature space $\phi(I)$ extracted with a convolutional stacked hourglass network significantly increases the quality of 3D human shapes. ReCHOR first extracts a feature embedding for each input image $\phi_{\{h,o,f\}}(I_{\{h,o,f\}})$. A first set of points $X_{h}$ is then projected onto $I_{f}$ and $I_{h}$ via perspective projection $\pi$ and linked to the corresponding features $\phi_{h}$ and $\phi_{f}$ to obtain pixel-alignment with the human: $\Phi_{\{h,f\}}^{h}=\phi_{\{h,f\}}(\pi(X_{h},I_{\{h,f\}}))$. Similarly, the pixel-aligned object features $\Phi_{\{o,f\}}^{o}=\phi_{\{o,f\}}(\pi(X_{o},I_{\{o,f\}}))$ are obtained by projecting a different set of points $X_{o}$ on $I_{o}$ and $I_{f}$ and indexing them with the corresponding features $\phi_{o}$ and $\phi_{f}$.
@@
Human-object pose prior features: Multiple 3D spatial configurations of humans and objects can project to the same 2D image, leading to difficulties in estimating their position in 3D space and depth-scale ambiguity. Since parametric model-based human-object reconstruction methods optimize 3D location, depth, and scale using geometric and spatial constraints, we leverage them to address depth-scale ambiguity and anchor the 3D spatial location. We condition our model on semantic features of the SMPL-H and object template estimated by the parametric model-based method, and we compute the neural representations of both human and object relative to the predicted SMPL-H center. The object position is defined relative to this center, to learn its 3D spatial relationship with the human. To define the human pose features, for a query point $x_{h}\in X_{h}$, we look for the closest point $x_{h}^{*}$ on SMPL-H, i.e. $x_{h}^{*}=\arg\min_{x_{h}^{p}}||x_{h}-x_{h}^{p}||$ where $x_{h}^{p}$ are points on the SMPL-H mesh. Human pose prior features $\sigma_{h}(x_{h})=[d_{h},v_{h},z_{h}]$ comprise three elements: a signed distance value $d_{h}$ between $x_{h}$ and $x_{h}^{*}$, a visibility label $v_{h}\in\{1,0\}$ where $v_{h}$ indicates if $x_{h}^{*}$ is visible in the image when SMPL-H and the object mesh is projected together, and a relative depth-aware feature $z_{h}=(x_{h}^{z}-z_{c})$ where $z_{c}$ is the depth of the SMPL-H center and $x_{h}^{z}$ is the depth of the query point $x_{h}$. Analogously, for a query point $x_{o}\in X_{o}$ and its closest point on the object mesh, $x_{o}^{*}$, the object pose prior features $\sigma_{o}(x_{o})=[d_{o},v_{o},z_{o}]$ are three elements: a signed distance value $d_{o}$ between $x_{o}$ and $x_{o}^{*}$, a visibility label $v_{o}\in\{1,0\}$ and a relative depth-aware feature $z_{o}=(x_{o}^{z}-z_{c})$ where $x_{o}^{z}$ is the depth of the query point $x_{o}$.
@@
Global-local context learning via transformer encoders: As shown in Sec. 4.2, simply concatenating the features of the human $\Phi_{h}^{h}$ and object $\Phi_{o}^{o}$ with those of the input image $\Phi_{f}^{\{h,o\}}$ deteriorates network performance, making it less robust to noise of the input image. Additionally, the network’s ability to integrate and reason about features across the entire image is limited by the receptive field of convolutional layers. To better understand the relationship between the human and the object in the space, we propose using two attention-based encoders, $A_{o}$ and $A_{h}$ that merge the feature extracted by the feature extractor module. An attention score is computed for the two paired images by assessing the compatibility between a query and its corresponding key, producing two feature embeddings: one for the human $\varphi_{h}=A_{h}(\Phi_{h}^{h},\Phi_{f}^{h},\sigma_{h}(x_{h}))$ and another for the object $\varphi_{o}=A_{o}(\Phi_{o}^{o},\Phi_{f}^{o},\sigma_{o}(x_{o}))$. The pose features $\sigma_{\{h,o\}}$ obtained in the previous module are also concatenated to integrate spatial information. Each feature $\varphi_{\{h,o\}}$ integrates the information from its corresponding local image $I_{\{h,o\}}$, combined with the global context of the input image $I_{f}$. A more comprehensive understanding of the scene is obtained by contextualizing the global and local information. As a result, the human feature $\varphi_{h}$ contains detailed information from the human image, as well as contextual information of the global scene from the input image. Similarly, the object feature $\varphi_{o}$ integrates scene-level information from $I_{f}$.